Please provide formula.
If a plane is flying at a constant of 550 mph for 6 hours and then returns in 5 hours what is the wind speed. The answer is 50 mph. But need formula
Thanks

Answer by mananth(16946)

6(550-x)=5(550+x)

3300-6x =2750+5x

11x=550
x=50